An investment of Ksh 100,000 in an account accumulated to Ksh 250,000 after 4 years.
a)State the accumulation factor A(0,4)
b)Find the simple annual interest rate which would give the accumulation factor in part (a)
1
Expert's answer
2020-07-27T18:54:31-0400

(a) . A(t1,t2) = T/C

where T = accumulated amount at the end of t2

C = the amount at the end of t1

given T = ksh250,000 , C = ksh100,000 and t1=0 , t2=4

A(0,4) = 250,000/100,000

A(0,4) = 2.5

(b) . for simple interest

A(t1,t2) = 1+ni where n = 4

A(0,4) = 1+ 4i

2.5 = 1+4i

i = (2.5 - 1)/4

i = 1.5 /4

i = 0.375

"\implies" 37.5%